# Inhomogeneous SSH models and the doubling of orthogonal polynomials

### v1

@@ 1 Introduction
The Su–Schrieffer–Heeger (SSH) model is a prominent example of such models, and was originally introduced to describe polyacetylene [14], where carbon atoms form a dimerized chain with alternating strong and weak bonds. Owing to this alternating bonding pattern, the model hosts symmetry-protected topological excitations and represents the simplest quantum many-body system exhibiting such features [15]. The SSH model is exactly solvable with free fermion techniques, but nonetheless describes experimentally realized topological phases [16], making it a paradigmatic framework for studying topological phase transitions in condensed matter systems.
@@
The paper is organized as follows: in Sec. 2 we show that the eigenvalues and eigenvectors of the SSH model can be related to the well-known Chebyshev polynomials and to their doubling. Section 3 extends the doubling procedure to any finite sequence of orthogonal polynomials, leading to new models that generalize the SSH model while remaining exactly solvable. Finally, in Secs. 4 and 5, we illustrate the general construction using two specific families of discrete polynomials from the ($q$-)Askey scheme: the Krawtchouk and $q$-Racah polynomials.
@@ 2 SSH model and doubling of Chebyshev polynomials
The SSH model is a one-dimensional free-fermion model with dimerized nearest-neighbor interactions. We consider a chain with an odd number of sites, labeled from 0 to $2N$ and endowed with open boundary conditions. The Hamiltonian $\mathcal{H}$ reads [14]
$\begin{split}\mathcal{H}=&\sum_{x=1}^{N}\big(t^{+}c_{2x-2}^{\dagger}c_{2x-1}+t^{-}c_{2x-1}^{\dagger}c_{2x}+\textrm{h.c.}\big)\\ =&\sum_{x,y=1}^{2N+1}H_{x,y}c_{x-1}^{\dagger}c_{y-1},\end{split}$ (2.1)
where $c_{x}^{(\dagger)}$ are the fermion operators satisfying the canonical anticommutation relations
$\{c_{x}^{\dagger},c_{y}\}=\delta_{x,y},\ \{c_{x},c_{y}\}=0=\{c_{x}^{\dagger},c_{y}^{\dagger}\},$ (2.2)
and $t^{\pm}$ are the dimerized interactions. For the standard SSH model, the hoping parameters read
@@
with $1\leqslant\delta\leqslant 1$. We represent the chain described by the Hamiltonian in Figure 1, where the alternating bounds are easily visualized.
@@
Because the full Hamiltonian $\mathcal{H}$ is quadratic in terms of the fermion operators, it suffices to diagonalize the $(2N+1)\times(2N+1)$ matrix $H$ to obtain the eigenvalues and eigenvectors of $\mathcal{H}$. This matrix $H$ is the single-particle Hamiltonian, and it reads
@@
In the following, we consider solvable models with dimerized couplings $t_{n}^{\pm}$ which depend on the position in the chain, thereby defining inhomogeneous SSH-like models. We represent a schematic inhomogeneous SSH model in Fig. 2. To begin this study, we show in the following subsection that the eigenvalues and the eigenvectors of the homogeneous SSH model can be obtained with the doubling [29] of Chebyshev polynomials.
[…]
Figure 2: Representation of an inhomogeneous SSH model.
2.1 Diagonalisation by Chebyshev polynomials
In this section we show that the Chebyshev polynomials can be used to diagonalize the matrix $H$ given in Eq. (2.4). To do this, let us recall some basic properties of the Chebyshev polynomials of the second kind, denoted as $U_{n}$ ($n=0,1,2,\dots$). They can be expressed in terms of hypergeometric functions as
@@ 2.1 Diagonalisation by Chebyshev polynomials
is an eigenvector of $H$ corresponding to the eigenvalue $x$, where $x$ is a root of the characteristic polynomial:
@@
It is important to emphasize that the parity of the number of sites plays a crucial role in obtainig analytical results. Indeed, when number of sites is equal to $2N$, the corresponding equation to determine the spectrum would be $Q_{2N}(x)=0.$ However the roots of this polynomial are not known explicitly in general.
@@ 2.2 SSH model with non-vanishing chemical potential
$0,\ x_{k}^{\pm}=\frac{\mu^{+}+\mu^{-}\pm\sqrt{(\mu^{+}-\mu^{-})^{2}+2(1+\delta^{2})+2(1-\delta^{2})\cos\left(\frac{k\pi}{N+1}\right)}}{2}\,,\quad(k=1,2,\dots,N)\,.$ (2.26)
@@ 3 Inhomogeneous SSH model from the doubling procedure
Inspired by the fact that the doubling procedure described in Remark 2.1 applies to any sequence of orthogonal polynomials, we can use the previous construction to obtain new models that generalize the SSH model and remain exactly solvable.
Instead of the Chebyshev polynomials $U_{n}$, let us consider a finite sequence of orthogonal polynomials $(P_{n})_{n=0}^{N-1}$ satisfying a three-term recurrence relation
$xP_{n}(x)=A_{n}P_{n+1}(x)-(A_{n}+C_{n})P_{n}(x)+C_{n}P_{n-1}(x)\,$, (3.1)
with the initial conditions $C_{0}=0,$ $A_{N-1}=0,$ $P_{0}(x)=1$. Assuming $A_{n}C_{n+1}>0$ (which we assume in the following), Favard’s theorem (see e.g. [33]) guarantees that the sequence $(P_{n})_{n}$ is orthogonal,
@@
with a certain weight function $\Omega(x)$ and norm ${\cal N}_{n}$. The sequence $(\lambda(x))_{x=0}^{N-1}$, appearing in the orthogonality relation, is called the grid associated to the polynomials $P_{n}$. The recurrence relation for n=N1n=N-1 is only valid for $x=\lambda(k)$ ($k=0,1,\dots,N-1$). We can change the normalization of the sequence $P_{n}$ to obtain a more convenient recurrence relation. Defining the sequence $(R_{n})_{n}$ as
@@
$Q_{2n}(x)=t^{+}_{n}R_{n}(\pi_{x})+t^{-}_{n-1}R_{n-1}(\pi_{x})\,$$\quad$ $(n=0,\dots,N)\,$, (3.5a)
$Q_{2n+1}(x)=xR_{n}(\pi_{x})\,$$\quad$ $(n=0,\dots,N-1)\,.$ (3.5b)
@@
where $\tau_{0},\tau_{2}$ are parameters that need to satisfy certain constraints (see below). The $Q_{n}(x)$ polynomials allow one to diagonalize a generalization of the SSH model, as per the following proposition.
Proposition 3.1.
The vector $\mathbb{Q}(x)$, whose components $Q_{0}(x),Q_{1}(x),\dots,Q_{2N}(x)$ are defined by (3.5), diagonalizes the $(2N+1)\times(2N+1)$ matrix
$\left(\begin{array}[]{c c c c c c c}0&t^{+}_{0}&\\ t^{+}_{0}&0&t^{-}_{0}\\ 0&t^{-}_{0}&0&t^{+}_{1}\\ &&&\ddots\\ &&&t^{+}_{N-1}&0&t^{-}_{N-1}\\ &&&&t^{-}_{N-1}&0\end{array}\right)\mathbb{Q}(x)=x\mathbb{Q}(x)\,$, (3.7)
with the eigenvalue $x$ chosen among
$0,\ x_{k}^{\pm}=\pm\sqrt{\frac{\lambda(k)-\tau_{0}}{\tau_{2}}}\,,\quad(k=0,1,\dots,N-1)\,$, (3.8)
if the following constraints are satisfied:
@@
$A_{n}+C_{n}+\tau_{0}=-\tau_{2}((t^{+}_{n})^{2}+(t^{-}_{n})^{2})\,.$ (3.9c)
Proof.
The components of the eigenvalue problem (3.7) read
$t_{n-1}^{-}Q_{2n-1}(x)+t_{n}^{+}Q_{2n+1}(x)=xQ_{2n}(x)\,,\quad(n=0,1,\dots,N)\,$, (3.10)
$t^{+}_{n}Q_{2n}(x)+t^{-}_{n}Q_{2n+2}(x)=xQ_{2n+1}(x)\,,\quad(n=0,1,\dots,N-1)\,.$ (3.11)
@@
The last equality is obtained using the three term recurrence relation for $R_{n}$. Using the constraints (3.9b) and (3.9c), one obtains
@@
which proves (3.11). Note that the three term recurrence relation used for (3.12) is valid only if $\pi_{x}=\lambda(k)$ ($k=0,1,\dots,N-1$) when n=N1n=N-1. Observing that (3.11) holds for $x=0$ and finding the roots $x$ of $\pi_{x}=\lambda(k)$, we obtain the eigenvalues (3.8). This concludes the proof of the proposition. ∎
The matrix in (3.7) can be interpreted as the Hamiltonian of an inhomogeneous SSH model which can be represented as in Fig. 2. Let us remark that it is important that $R_{n}$ be a finite sequence for $n=0,1,\dots,N-1$ such that the spectrum is analytically computable.

### v2

@@ 1 Introduction
The Su–Schrieffer–Heeger (SSH) model is a prominent example of such models, and was originally introduced to describe polyacetylene [14], where carbon atoms form a dimerized chain with alternating strong and weak bonds. Owing to this alternating bonding pattern, the model hosts symmetry-protected topological excitations and represents the simplest quantum many-body system exhibiting such features [15]. The SSH model is exactly solvable with free-fermion techniques, but nonetheless describes experimentally realized topological phases [16, 17, 18], making it a paradigmatic framework for studying topological phase transitions in condensed matter systems.
@@
While the inhomogeneous couplings generated by orthogonal-polynomial constructions do not aim to describe a specific crystalline material, they naturally correspond to engineered hopping profiles. Such spatially modulated SSH-type models are experimentally relevant in platforms where couplings can be controlled with high precision, including cold atoms in optical lattices, photonic or acoustic lattices, and other synthetic quantum systems [16, 17, 18]. In this sense, these models provide analytically tractable realizations of inhomogeneous topological chains rather than microscopic, material-specific descriptions.
Main results.
The Hamiltonian of the inhomogeneous models studied in this paper has the form
$\mathcal{H}=\sum_{n=1}^{N}\big(t_{n-1}^{+}c_{2n-2}^{\dagger}c_{2n-1}+t_{n-1}^{-}c_{2n-1}^{\dagger}c_{2n}+\textrm{h.c.}\big),$ (1.1)
where $t^{\pm}_{n}$ are parameters characterizing the strength of the interaction between neighboring sites, and $c_{n}^{(\dagger)}$ are the fermionic operators satisfying the canonical anticommutation relations
$\{c_{m}^{\dagger},c_{n}\}=\delta_{m,n},\ \{c_{m},c_{n}\}=0=\{c_{m}^{\dagger},c_{n}^{\dagger}\}\;,\quad(m,n=0,1,\dots,N).$ (1.2)
We represent a schematic inhomogeneous SSH model in Fig. 1.
[…]
Figure 1: Representation of an inhomogeneous SSH model.
The doubling procedure allows us to obtain the coupling constants $t^{\pm}_{n}$ such that the spectrum and the eigenvectors can be computed exactly. The different models obtained, with the names of the polynomials used in the doubling procedure, are summarized in the Table 1 given below.
@@
The paper is organized as follows: in Sec. 2 we show that the eigenvalues and eigenvectors of the standard SSH model can be related to the well-known Chebyshev polynomials and to their doubling. Section 3 extends the doubling procedure to any finite sequence of orthogonal polynomials, leading to new models that generalize the SSH model while remaining exactly solvable. Finally, in Secs. 4 and 5, we illustrate the general construction using two specific families of discrete polynomials from the ($q$-)Askey scheme: the Krawtchouk and $q$-Racah polynomials.
@@ 2 SSH model and doubling of Chebyshev polynomials
This section is devoted to study the standard SSH model, where the coupling constant are constant along the chain: $t^{+}_{n}=t^{+}$ and $t^{-}_{n}=t^{-}$. We consider a chain with an odd number of sites, labeled from 0 to $2N$ and endowed with open boundary conditions. The Hamiltonian $\mathcal{H}$ for the homogeneous SSH model reads [14]
$\begin{split}\mathcal{H}=&\sum_{n=1}^{N}\big(t^{+}c_{2n-2}^{\dagger}c_{2n-1}+t^{-}c_{2n-1}^{\dagger}c_{2n}+\textrm{h.c.}\big)\\ =&\sum_{m,n=1}^{2N+1}H_{m,n}c_{m-1}^{\dagger}c_{n-1},\end{split}$ (2.1)
where $c_{n}^{(\dagger)}$ are the fermionic operators satisfying the canonical anticommutation relations given in Eq. (1.2). The dimerized interactions $t^{\pm}$ read
@@
with $1\leqslant\delta\leqslant 1$. We represent the chain described by this Hamiltonian in Figure 2, where the alternating bounds are easily visualized.
@@
Because the full Hamiltonian $\mathcal{H}$ is quadratic in terms of the fermionic operators, it is well-known that it suffices to diagonalize a $(2N+1)\times(2N+1)$ matrix $H$ to obtain the eigenvalues and eigenvectors of $\mathcal{H}$. This matrix $H$ is the single-particle Hamiltonian, and it reads
@@
2.1 Diagonalization by Chebyshev polynomials
The diagonalization of this homogeneous SSH model is standard, see e.g. [35, 36, 37]. However, to obtain exactly solvable inhomogeneous SSH model, we start by showing that the eigenvalues and eigenvectors of the homogeneous SSH model can be obtained using the doubling property of Chebyshev polynomials [31]. This offers a new mathematical framework for the standard diagonalization of the homogeneous model, and it extends naturally to inhomogeneous cases.
Let us recall some basic properties of the Chebyshev polynomials of the second kind, denoted as $U_{n}$ ($n=0,1,2,\dots$). They can be expressed in terms of hypergeometric functions as
@@ 2.1 Diagonalization by Chebyshev polynomials
is an eigenvector of $H$ (2.3) corresponding to the eigenvalue $x$, where $x$ is a root of the characteristic polynomial:
@@
The eigenvector associated to the eigenvalue $x=0$ is known as a zero mode, which is a fundamental feature of topological insulators [37, 15]. In the standard SSH model with odd number of sites, the zero mode is localized at the right (resp. left) boundary of the chain for $\delta>0$ (resp. $\delta<0$). These two situations correspond to distinct topological phases of the model. For $\delta=0$, the model is instead critical, or gapless.
It is important to emphasize that the parity of the number of sites plays a crucial role in obtaining analytical results. Indeed, when the number of sites is even, equal to $2N$, the corresponding equation to determine the spectrum would be $Q_{2N}(x)=0.$ However the roots of this polynomial are not known explicitly in general.
@@ 2.2 SSH model with non-vanishing chemical potential
$\mu^{+},\ x_{k}^{\pm}=\frac{\mu^{+}+\mu^{-}\pm\sqrt{(\mu^{+}-\mu^{-})^{2}+2(1+\delta^{2})+2(1-\delta^{2})\cos\left(\frac{k\pi}{N+1}\right)}}{2}\,,\quad(k=1,2,\dots,N)\,.$ (2.25)
@@ 3 Inhomogeneous SSH model from the doubling procedure
Inspired by the fact that the doubling procedure described in Remark 2.1 applies to any sequence of orthogonal polynomials, we can use the previous construction to obtain inhomogeneous models (1.1) that generalize the SSH model and remain exactly solvable. In this case, the tridiagonal matrix to diagonalize reads as
$H^{(inh)}=\left(\begin{array}[]{c c c c c c c}0&t^{+}_{0}&\\ t^{+}_{0}&0&t^{-}_{0}\\ 0&t^{-}_{0}&0&t^{+}_{1}\\ &&&\ddots\\ &&&t^{+}_{N-1}&0&t^{-}_{N-1}\\ &&&&t^{-}_{N-1}&0\end{array}\right)\,.$ (3.1)
We first describes explicitly the zero mode for this matrix. Next, we recall well-known properties of orthogonal polynomials necessary for the following construction. Using this framework, we propose an ansatz, based on the doubling procedure, for the eigenvectors of the inhomogeneous models. Finally, we establish the parameter constraints required to ensure the existence of these eigenvectors.
3.1 Zero mode
The matrix in (3.1) can be interpreted as the single-particle Hamiltonian of an inhomogeneous SSH model which is represented in Fig. 1. The existence of a zero mode is a direct consequence of the chiral symmetry of the model (see e.g. [37, 15]). One can check that the vector $\textbf{Q}^{(0)}$, whose components $Q_{0}^{(0)},Q_{1}^{(0)},\dots,Q_{2N}^{(0)}$ are defined by
$Q_{2n+1}^{(0)}=0\,,\quad(n=0,1,\dots,N-1),\qquad Q_{2n}^{(0)}=(-1)^{n}\prod_{i=0}^{n-1}\frac{t_{i}^{+}}{t_{i}^{-}}\,,\quad(n=0,1,\dots,N)\,$, (3.2)
is an eigenvector of $H^{(inh)}$ with eigenvalue 0. We will describe the other modes in Subsec. 3.3. The existence of this zero mode is the main reason to focus the study on models with odd numbers of sites. We will also investigate a model with an even number of sites in Subsec. 5.3.
For the homogeneous SSH chain with $t^{+}_{n}=t^{+}$ and $t^{-}_{n}=t^{-}$, related to Chebyshev polynomials, the zero mode is localized at the end of the chain, as discussed in Sec. 2.1. From Eq. (3.2), it is direct to conclude that the zero mode is localized on the last site of the chain for $t^{+}>t^{-}$, whereas it is localized on the first site for $t^{+}<t^{-}$. In the inhomogeneous case, we focus on functions $t^{+}_{n}$ which monotonically decrease along the chain, whereas functions $t^{-}_{n}$ monotonically increase. In this case, unlike for the homogeneous SSH chain, we expect the zero mode to be localized in regions of space where $t_{n}^{+}\sim t_{n}^{-}$, thereby separating regions where $t_{n}^{+}>t_{n}^{-}$ and $t_{n}^{+}<t_{n}^{-}$, which correspond to different topological phases. This behavior is again a direct consequence of Eq. (3.2). Of course, the precise extension of the zero mode and the validity of this conjecture need to be verified in specific examples, which will be addressed in a forthcoming publication [38].
3.2 Orthogonal polynomials
Hypergeometric orthogonal polynomials, along with their $q$-analogs [39, 33, 40, 34], form a central class of special functions. These families of polynomials are systematically organized in the ($q$-)Askey scheme and are characterized by a three-term recurrence relation in view of Favard’s theorem. Moreover, they exhibit a bispectral property: they satisfy both a recurrence relation in the degree and a second-order differential or ($q$-)difference equation in the variable, which justifies their classification as classical polynomials. Understanding the detailed properties of these functions is valuable, as it enables their effective application in diverse mathematical and physical contexts. The families of orthogonal polynomials in the ($q$-)Askey scheme are defined in terms of ($q$-)hypergeometric functions (1.5). For our purposes, we concentrate on the finite families within the ($q$-)Askey scheme, the most general of which are the ($q$-)Racah polynomials, and we consider only the properties necessary for the subsequent analysis.
We denote by $P_{n}(x)=P_{n}(x;{\bm{\rho}})$, with $n=0,1,\dots,N-1$ and $N$ a non-negative integer, the polynomials of a given finite family. Here, ${\bm{\rho}}=(\rho_{1},\rho_{2},\dots)$ represents the set of parameters that define the family, including the integer $N$. These polynomials satisfy a three-term recurrence relation
$xP_{n}(x)=A_{n}P_{n+1}(x)-(A_{n}+C_{n})P_{n}(x)+C_{n}P_{n-1}(x)\,,\quad(n=0,1,\dots,N-1)$ (3.3)
with the initial conditions $C_{0}=0,$ $A_{N-1}=0,$ $P_{0}(x)=1$. The previous relations for $n=0,1,\dots,N-2$ allows us to define recursively the polynomials $P_{1}$, $P_{2},\dots,P_{N-1}$. The last relation for n=N1n=N-1 reads $xP_{N-1}(x)=C_{N-1}(P_{N-2}(x)-P_{N-1}(x))$ and it provides a constraint for $N$ possible values of $x$ found as the roots of this equation. We denote these values $(\lambda(x))_{x=0}^{N-1}$ and call it the grid associated to the polynomials $P_{n}$. Assuming $A_{n}C_{n+1}>0$ (which we assume in the following), Favard’s theorem (see e.g. [40]) guarantees that the sequence $(P_{n})_{n}$ is orthogonal,
@@ 3.2 Orthogonal polynomials
with a certain weight function $\Omega(x)$ and norm ${\cal N}_{n}$. The finite families possess several important additional properties. One of these is known as Wilson duality: for any polynomials $P_{n}(x;{\bm{\rho}})$, there exist polynomials $\mathcal{P}_{i}(x;{\bm{\rho}}^{\prime})$ such that
$\mathcal{P}_{n}(x;{\bm{\rho}}^{\prime})=P_{x}(n;{\bm{\rho}}).$ (3.5)
Additionally, they satisfy contiguity relations, that is, equations of the form
$\lambda^{+}_{x;{\bm{\rho}}}\,R_{i}(x;{\bm{\rho}})$ $=\sum_{\epsilon\in\mathscr{S}}\Phi^{\epsilon,+}_{i}\,R_{i+\epsilon}({\overline{x}};\overline{{\bm{\rho}}})\,$, (3.6)
where $\overline{x}$} denotes a shift of $x$, namely ${\overline{x}}=x+\eta$ with $\eta\in\{0,+1,-1\}$, $\overline{{\bm{\rho}}}$ is a set of modified parameters, and $\mathscr{S}$ is a finite set of integers. In [41], contiguity relations for the finite families of the Askey scheme have been classified, with $\mathscr{S}$ taken to be one of the sets $\{0,-1\}$, $\{0,-1,1\}$, or $\{0,-1,-2\}$. We note that some of these contiguity relations are commonly referred to as (forward or backward) shift relations [34].
3.3 Solving the inhomogeneous SSH model
In what follows, we change the normalization of the sequence $P_{n}$ in order to obtain a more convenient recurrence relation. Defining the sequence $(R_{n})_{n}$ as
@@ 3.3 Solving the inhomogeneous SSH model
$Q_{2n}(x)=t^{+}_{n}R_{n}(\pi_{x})+t^{-}_{n-1}R_{n-1}(\pi_{x})\,$, $(n=0,\dots,N)\,$, (3.9a)
$Q_{2n+1}(x)=xR_{n}(\pi_{x})\,$, $(n=0,\dots,N-1)\,.$ (3.9b)
@@
where $\tau_{0},\tau_{2}$ are parameters that need to satisfy certain constraints (see below). The $Q_{n}(x)$ polynomials allow one to diagonalize the matrix $H^{(inh)}$ defined in Eq. (3.1) associated to an inhomogeneous SSH model, as we show in the following paragraph.
Inhomogeneous SSH models.
If the coupling constants $t^{\pm}_{n}$ of the models are given by the following relations in terms of the recurrence relation coefficients $A_{n}$, $C_{n}$ of a sequence of polynomials:
@@
$A_{n}+C_{n}+\tau_{0}=-\tau_{2}((t^{+}_{n})^{2}+(t^{-}_{n})^{2})\,$, (3.11c)
then the vector $\mathbb{Q}(x)$, whose components $Q_{0}(x),Q_{1}(x),\dots,Q_{2N}(x)$ are defined by the previous ansatz (3.9), diagonalizes the $(2N+1)\times(2N+1)$ matrix $H^{(inh)}$ (3.1),
$H^{(inh)}\mathbb{Q}(x)=x\mathbb{Q}(x)\,.$ (3.12)
The associated eigenvalues $x$ are given by
$x_{k}^{\pm}=\pm\sqrt{\frac{\lambda(k)-\tau_{0}}{\tau_{2}}}\,,\quad(k=0,1,\dots,N-1)\,$, (3.13)
where $\lambda(k)$ is the grid associated to the orthogonal polynomials $(P_{n}(x))_{n=0}^{N-1}$. We recall that $H^{(inh)}$ has also a vanishing eigenvalue with the eigenvector given by (3.2).
The remaining part of this paragraph is devoted to the proof of these results. The components of the eigenvalue problem (3.12) read
$t_{n-1}^{-}Q_{2n-1}(x)+t_{n}^{+}Q_{2n+1}(x)=xQ_{2n}(x)\,$, $(n=0,1,\dots,N)\,$, (3.14)
$t^{+}_{n}Q_{2n}(x)+t^{-}_{n}Q_{2n+2}(x)=xQ_{2n+1}(x)\,$, $(n=0,1,\dots,N-1)\,.$ (3.15)
@@
The last equality is obtained using the three-term recurrence relation for $R_{n}$. Using the constraints (3.11b) and (3.11c), one obtains
@@
which proves (3.15). As mentioned after (3.3), the three-term recurrence relation used for (3.16) is valid only if $\pi_{x}$ is on the grid of the associated polynomials i.e. $\pi_{x}=\lambda(k)$ ($k=0,1,\dots,N-1$). The eigenvalues (3.13) are then obtained when solving $\pi_{x}=\lambda_{k}$.
Finding the eigenvalues.
We already showed that $x=0$ is an eigenvalue with eigenvector $\textbf{Q}^{(0)}$ defined by (3.2). The vectors $\textbf{Q}(x)$ are eigenvectors if $x=x_{k}$ for some $k=0,\cdots,N-1$. Indeed, equation (3.14) is satisfied for all $x$ and all $n=0\cdots N$. The only issue appears in equation (3.15) for n=N1n=N-1. In fact, we had to use the recurrence relation for the family $R_{n}$ which is valid only if $\pi_{x}$ is on the grid $(\lambda(k))$. Solving $\pi_{x}=\lambda(k)$ gives the eigenvalues $x_{k}^{\pm}$ (3.13).
Trivial solution.